How do I make a negative exponent that is in the denominator positive? Do I just move the coefficient to the numerator?
Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

How do I make a negative exponent that is in the denominator positive? Do I just move the coefficient to the numerator?
-----
That is what it looks like.
1/a^-1 = a
----
Why???
1/a^-1 = 1/(1/a) = [a/a][1/(1/a)] = a/1 = a
================================================
Cheers,
Stan H.

Answer by solver91311(24713)



You move the variable and its exponent but not the coefficient if there is one to the numerator and drop the minus sign on the exponent.



But:



John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism